Im doing my homework and i've been on the same question for like 10 minutes
The question is:
k over 5 - 1 =7
And i cant figure out how to do the work one this problem. Thankyou!

Answer by checkley79(3341)

k over 5 - 1 =7
IF YOU MEAN:
K/(5-1)=7
K/4=7 CROSS MULTIPLY
K=4*7
K=28 ANS.
IF YOU MEAN:
K/5-1=7
K/5=7+1
K/5=8 CROSS MULTIPLY.
K=5*8
K=40 ANS.